# Investigating the Robustness of Extreme Precipitation Super-Resolution Across Climates

### v2

@@ 3.2 At Which Level of Coarsening Does Super-Resolution Fail?
For clarity of exposition, we assume the existence of two regular grids, $\mathcal{G}_{m}^{\delta_{1}}$ and $\mathcal{G}_{m}^{\delta_{2}}$, as introduced in Section 3.1, with $\delta_{1}>\delta_{2}>0$, and where $\delta_{1}$ is an integer multiple of $\delta_{2}$. While this assumption simplifies the presentation, our approach naturally extends to more general settings involving arbitrary grids.
A natural approach for addressing the question posed in this section is to examine the spatial correlation function of the field of interest (e.g., precipitation intensity), denoted by $X$. This function captures how similar the values of the field are at different sites, depending on their spatial separation—providing insight into the scale and structure of spatial patterns, such as whether events are localized or spread out over larger regions. In particular, it allows us to assess whether the value of the field at a distant site still carries meaningful information about the value at a reference point, which is crucial for identifying the resolution below which super-resolution techniques may no longer yield significant benefits. However, in most cases, the field $X$ is not stationary, and we therefore standardize it so that it has zero mean and unit variance at each grid point, resulting in a field denoted by $\tilde{X}$. In practice, we compute, for each $\bm{s}\in\mathcal{G}_{m}^{\delta_{2}}$, the temporal mean and standard deviation
@@
$\tilde{X}(s)\overset{\mathrm{def}}{=}\frac{X(\bm{s})-\bar{X}(\bm{s})}{\hat{\mathrm{std}}_{X}^{\text{(t)}}(\bm{s})},\quad\bm{s}\in\mathcal{G}_{m}^{\delta_{2}},$ (5)
where $N$ denotes the number of time points (equal to $33$ in our application) and $X_{i}$ is the field $X$ observed at time $i$. This standardization has the advantage that the resulting field $\tilde{X}$ can reasonably be treated as second-order stationary, meaning it has a constant mean across space and a covariance structure that depends only on the relative displacement between sites, not their absolute positions.
We denote by $\rho_{\tilde{X}}(\bm{r})$, for $\bm{r}\in\mathcal{G}_{m}^{\delta_{2}}$, the spatial correlation function of the transformed field. Under the assumption of isotropy, which we adopt here, this function depends only on the distance between sites, and is thus written $\rho_{\tilde{X}}(r)$. We define
$B_{r}\overset{\mathrm{def}}{=}\left\{(\bm{s}_{1},\bm{s}_{2})\in\mathcal{G}_{m}^{\delta_{2}}:d(\bm{s}_{1},\bm{s}_{2})\in\left[r-\frac{\Delta r}{2},\,r+\frac{\Delta r}{2}\right]\right\}$ (6)
@@ 3.3 Quantifying Generalizability and Robustness across Climates
The overall goal of this section is to (i) formalize the notions of generalization and robustness abilities of a model in a climate change context; (ii) propose a concrete solution tailored to estimating the quantiles of fine-resolution distributions. To do this, we consider three strategies: (i) a model trained on present-day data and applied to present conditions; (ii) a model trained on present-day data and applied to future conditions; and (iii) a model both trained and evaluated on future data.
Let $y$ denote an arbitrary characteristic of the distribution under consideration (e.g., the mean, a quantile, etc.) at a given grid point. Throughout this section, to maintain notational simplicity, we omit explicit reference to grid-point specificity and use lowercase letters even for random variables. Let $y_{\mathrm{P}}^{\mathrm{P}}$ be the predicted quantity for the present climate using a model trained on present-day data, $y_{\mathrm{F}}^{\mathrm{P}}$ be the predicted quantity for the future climate using a model trained on present-day data, and $y_{\mathrm{F}}^{\mathrm{F}}$ be the predicted quantity for the future climate using a model trained on future data. Moreover, let $y^{\mathrm{P}}$ and $y^{\mathrm{F}}$ denote the observed (true) quantities in the present and future climates, respectively. Let $\ell$ be a generic point-wise loss function that measures the discrepancy between predicted and true values, and let $\mathbb{E}$ denote the expectation operator with respect to the spatial dimension. The use of this expectation enables the integration of pointwise losses over the entire grid. A concrete example of loss function is the squared distance, defined by $\ell(y,\hat{y})=(y-\hat{y})^{2}$.
We can now define the generalization gap (that can also be called extrapolation gap in our context), as
$\mathrm{Generalization\ Gap}\overset{\mathrm{def}}{=}\mathbb{E}\left[\ell(y^{\text{P}},y_{\text{P}}^{\text{P}})\right]-\mathbb{E}\left[\ell(y^{\text{F}},y_{\text{F}}^{\text{P}})\right],$ (11)
which quantifies the performance gap between a model trained and evaluated on present-day data versus the same model evaluated on future data. It captures both the inadequacy of parameters learned under current climate conditions when applied to future climates, and the impact of shifts in the model’s covariates over time. Thus, this generalization gap can be challenging to interpret, as a high value—indicating poor performance under future climate conditions—may primarily result from shifts in covariates, reflecting fundamentally different environmental conditions, rather than a failure of the model itself.
@@
$\mathrm{Robustness\ Gap}\overset{\mathrm{def}}{=}\mathbb{E}\left[\ell(y^{\text{F}},y_{\text{F}}^{\text{P}})\right]-\mathbb{E}\left[\ell(y^{\text{F}},y_{\text{F}}^{\text{F}})\right].$ (12)
This metric compares two models evaluated on the same future data, effectively removing the influence of covariate shifts. It quantifies how well a model trained on present-day data performs under future conditions, using a model trained on future-data as reference, thereby serving as an indicator of its transferability across climates. A low value indicates that the model parameters are robust to climate shifts, whereas a high one suggests that a model trained on present-day data may not be directly applicable under future climate conditions.
@@
Let $q_{\text{F},\alpha}^{\text{P}}$ denote the predicted quantile for the future climate obtained using a model trained on present-day data, and $q_{\text{F},\alpha}^{\text{F}}$ the predicted quantile obtained using a model trained on future data. The observed (true) quantiles in the present and future climate are denoted by $q_{\alpha}^{\text{P}}$ and $q_{\alpha}^{\text{F}}$, respectively. In this context, the robustness gap introduced above becomes
$\mathrm{Robustness\ Gap}=\mathbb{E}\left[\ell(q_{\alpha}^{\text{F}},q_{\text{F},\alpha}^{\text{P}})-\ell(q_{\alpha}^{\text{F}},q_{\text{F},\alpha}^{\text{F}})\right].$ (13)
@@
We now introduce two interpretable quantities that will naturally appear in the decomposition of the pointwise robustness gap (see below). Let $\varepsilon_{\text{F},\alpha}=q_{\text{F},\alpha}^{\text{F}}-q_{\alpha}^{\text{F}}$, which is the fit bias of the model trained on the future climate for the quantile at level $\alpha$, and $\Delta_{\alpha}=q_{\text{F},\alpha}^{\text{P}}-q_{\text{F},\alpha}^{\text{F}}$, as it quantifies the sensitivity of the predicted quantile to the choice of training data (whether from the present-day or future climate). We expect $\varepsilon_{F,\alpha}$ to be small if the model is adequate and has been properly trained. With these notations, it is straightforward to obtain that
$q_{\text{F},\alpha}^{\text{F}}=q_{\alpha}^{\text{F}}+\varepsilon_{\text{F},\alpha},\quad q_{\text{F},\alpha}^{\text{P}}=q_{\alpha}^{\text{F}}+\varepsilon_{\text{F},\alpha}+\Delta_{\alpha},$ (15)
and the term within the expectation in (13) is
$\quad\ \ \mathrm{Pointwise\ Robustness\ Gap}$
$\overset{\mathrm{def}}{=}\ell_{\alpha}\left(q_{\alpha}^{\text{F}},q^{\text{P}}_{\text{F},\alpha}\right)-\ell_{\alpha}\left(q_{\alpha}^{\text{F}},q^{\text{F}}_{\text{F},\alpha}\right)$
$=\left(\alpha-\mathbb{I}_{\{q_{\alpha}^{\text{F}}<q^{\text{P}}_{\text{F},\alpha}\}}\right)\left(q_{\alpha}^{\text{F}}-q^{\text{P}}_{\text{F},\alpha}\right)-\left(\alpha-\mathbb{I}_{\{q_{\alpha}^{\text{F}}<q^{\text{F}}_{\text{F},\alpha}\}}\right)\left(q_{\alpha}^{\text{F}}-q^{\text{F}}_{\text{F},\alpha}\right)$
$=-\alpha\left(q^{\text{P}}_{\text{F},\alpha}-q^{\text{F}}_{\text{F},\alpha}\right)-\mathbb{I}_{\{q_{\alpha}^{\text{F}}<q^{\text{P}}_{\text{F},\alpha}\}}\left(q_{\alpha}^{\text{F}}-q^{\text{P}}_{\text{F},\alpha}\right)+\mathbb{I}_{\{y^{\text{F}}<q^{\text{F}}_{\text{F},\alpha}\}}\left(q_{\alpha}^{\text{F}}-q^{\text{F}}_{\text{F},\alpha}\right)$
$=-\alpha\Delta_{\alpha}+\mathbb{I}_{\{\varepsilon_{\text{F},\alpha}+\Delta_{\alpha}>0\}}\left(\varepsilon_{\text{F},\alpha}+\Delta_{\alpha}\right)-\mathbb{I}_{\{\varepsilon_{\text{F},\alpha}>0\}}\varepsilon_{\text{F},\alpha}$
@@
$\mathrm{Pointwise\ Robustness\ Gap}=\begin{cases}(1-\alpha)\Delta_{\alpha},&\Delta_{\alpha}>0,\\ \alpha|\Delta_{\alpha}|,&\Delta_{\alpha}<0,\end{cases}$ (17)
@@
When incorporated into (3.3), this decomposition establishes a general framework for analyzing the contribution of individual parameter shifts to the pointwise robustness gap. It maintains the generality of the expression, ensuring applicability across a wide range of parametric distributions. Overall, our framework provides a comprehensive understanding of the robustness gap—clarifying not only when the super-resolution model generalizes effectively, when its performance begins to degrade, and when it ultimately fails, but also uncovering the underlying factors driving these behaviors.
@@ 4.1 Generalized Extreme-Value Distribution
To characterize extreme hourly precipitation, we adopt the Generalized Extreme-Value framework from Extreme-Value Theory (EVT), a branch of statistics focused on modeling the behavior of distribution tails. Unlike classical methods that describe central tendencies such as means and variances, EVT provides asymptotically justified models for block maxima or threshold exceedances, making it particularly well suited for assessing the risk of rare, high-impact events like extreme hourly rainfall (Coles, 2001).
@@ 4.2 Incorporating Features in the GEV Parameters
We now specify the form of the function $f$ introduced in (4), adapted to our specific setting in which the underlying parametric distribution is the GEV distribution. To flexibly capture how the features (parameters of the GEV distributions at coarse-resolution and auxiliary features) influence each parameter of the GEV distribution, we use the Vector Generalized Additive Model (VGAM) framework (Yee, 2015). The VGAM extends the familiar Generalized Additive Model (GAM) by allowing for multiple response variables. In our case the vector of response variables is composed of the GEV parameters.
@@
We typically model $\mu$ and $\log\sigma$ as smooth functions of the features but, for stability and identifiability, take $\xi$ to be constant (i.e., $f_{\mu,k}\equiv 0$ for all $k$). Indeed the estimation of $\xi$ is notoriously imprecise even with large samples of block maxima and allowing $\xi$ to vary with covariates often yields unstable fits (Coles, 2001; Davison and Smith, 1990). Moreover, extreme events are inherently rare, so there is typically insufficient data within each covariate “slice” to support a reliable smooth trend in $\xi$, and attempts to do so can lead to over‐parameterization and degraded predictive performance (Eastoe and Tawn, 2009). Finally, empirical comparisons indicate that introducing time‐ or covariate‐dependence in the location (and sometimes scale) parameters captures the bulk of observed non‐stationarity, while varying $\xi$ provides only marginal improvements at the cost of substantially increased uncertainty (Katz et al., 2002; Coles, 2001).

### v3

@@ 3.2 At Which Level of Coarsening Does Super-Resolution Fail?
For clarity of exposition, we assume the existence of two regular grids, $\mathcal{G}^{\delta_{1}}$ and $\mathcal{G}^{\delta_{2}}$, as introduced in Section 3.1, with $\delta_{1}>\delta_{2}>0$, and where $\delta_{1}$ is an integer multiple of $\delta_{2}$. While this assumption simplifies the presentation, our approach naturally extends to more general settings involving arbitrary grids.
A natural approach for addressing the question posed in this section is to examine the spatial correlation function of the field of interest (e.g., precipitation intensity), denoted by $X$. This function captures how similar the values of the field are at different sites, depending on their spatial separation—providing insight into the scale and structure of spatial patterns, such as whether events are localized or spread out over larger regions. In particular, it allows us to assess whether the value of the field at a distant site still carries meaningful information about the value at a reference point, which is crucial for identifying the resolution below which super-resolution techniques may no longer yield significant benefits. However, in most cases, the field $X$ is not stationary, and we therefore standardize it so that it has zero mean and unit variance at each grid point, resulting in a field denoted by $\tilde{X}$. In practice, we compute, for each $\bm{s}\in\mathcal{G}^{\delta_{2}}$, the temporal mean and standard deviation
@@
$\tilde{X}(\bm{s})\overset{\mathrm{def}}{=}\frac{X(\bm{s})-\bar{X}(\bm{s})}{\hat{\mathrm{std}}_{X}^{\text{(t)}}(\bm{s})},\quad\bm{s}\in\mathcal{G}^{\delta_{2}},$ (5)
where $N$ denotes the number of time points (equal to $33$ in our application) and $X_{i}$ is the field $X$ observed at time $i$. This standardization provides an approximate treatment toward second-order stationarity, as the resulting field $\tilde{X}$ has constant mean and variance across space. However, this is not a full stationarization: correlations may still depend on absolute gridpoint positions rather than solely on relative displacement.
We denote by $\rho_{\tilde{X}}(\bm{r})$, for $\bm{r}\in\mathcal{G}^{\delta_{2}}$, the spatial correlation function of the transformed field. Under the assumption of isotropy, which we adopt here, this function depends only on the distance between sites, and is thus written $\rho_{\tilde{X}}(r)$. We define
$B_{r}\overset{\mathrm{def}}{=}\left\{(\bm{s}_{1},\bm{s}_{2})\in\mathcal{G}^{\delta_{2}}:d(\bm{s}_{1},\bm{s}_{2})\in\left[r-\frac{\Delta r}{2},\,r+\frac{\Delta r}{2}\right]\right\}$ (6)
@@
While we focus here on the spatial correlation function as a measure of dependence, it is important to acknowledge that correlation only captures linear relationships and may not fully represent the complex dependencies associated with extremes. However, by construction, the monthly maxima at lower resolution are obtained through (i) mean pooling of the original hourly data at 2.2 km resolution and (ii) taking the monthly maxima of these coarse-grained observations (see Section 2.2 for details). Consequently, the monthly maxima at the lower resolution still involve values at the 2.2 km scale that are not necessarily extreme, which is why we consider the spatial correlation function an acceptable metric in this context. Nevertheless, the proposed methodology is general and could be applied using alternative dependence measures, such as the extremal coefficient or other metrics tailored to extreme events.
@@ 3.3 Quantifying Generalizability and Robustness across Climates
The overall goal of this section is to (i) formalize the notions of generalization and robustness abilities of a model in a climate change context; (ii) propose a concrete solution tailored to estimate the quantiles of fine-resolution distributions. To do this, we consider three strategies: (i) a model trained on present-day data and applied to present conditions; (ii) a model trained on present-day data and applied to future conditions; and (iii) a model both trained and evaluated on future data.
Throughout this section, for notational simplicity, variables may refer either to observations at a specific grid point or to the entire field across all grid points, as should be clear from the context. Let $y$ denote an arbitrary characteristic of the distribution under consideration (e.g., the mean or a quantile). Let $y_{\mathrm{P}}^{\mathrm{P}}$ be the predicted quantity for the present climate using a model trained on present-day data, $y_{\mathrm{F}}^{\mathrm{P}}$ be the predicted quantity for the future climate using a model trained on present-day data, and $y_{\mathrm{F}}^{\mathrm{F}}$ be the predicted quantity for the future climate using a model trained on future data. Moreover, let $y_{\mathrm{P}}$ and $y_{\mathrm{F}}$ denote the observed (true) quantities in the present and future climates, respectively. Let $\ell$ be a generic point-wise loss function that measures the discrepancy between predicted and true values. A common example is the squared error, defined by $\ell(y,\hat{y})=(y-\hat{y})^{2}$.
3.3.1 Robustness Gap and Normalization Strategies
We can now define the generalization gap (also referred to as extrapolation gap in our context), as
$\mathrm{GG}\overset{\mathrm{def}}{=}\left\langle\ell\left(y_{\text{P}},y_{\text{P}}^{\text{P}}\right)\right\rangle-\left\langle\ell\left(y_{\text{F}},y_{\text{F}}^{\text{P}}\right)\right\rangle,$ (11)
which quantifies the performance gap between a model trained and evaluated on present-day data versus the same model evaluated on future data. We recall that the notation $\langle\cdot\rangle$ denotes the spatial mean, i.e., the average of the field over all grid points. The use of this operator enables the integration of pointwise losses over the entire grid. It captures both the inadequacy of parameters learned under current climate conditions when applied to future climates, and the impact of shifts in the model’s covariates over time. Thus, this generalization gap can be challenging to interpret, as a high value—indicating poor performance under future climate conditions—may primarily result from shifts in covariates, reflecting fundamentally different environmental conditions, rather than a failure of the model itself.
@@ 3.3.1 Robustness Gap and Normalization Strategies
$\mathrm{RG}\overset{\mathrm{def}}{=}\left\langle\ell\left(y_{\text{F}},y_{\text{F}}^{\text{P}}\right)\right\rangle-\left\langle\ell\left(y_{\text{F}},y_{\text{F}}^{\text{F}}\right)\right\rangle.$ (12)
This metric compares two models evaluated on the same future data, effectively removing the influence of covariate shifts. It quantifies how well a model trained on present-day data performs under future conditions, using a model trained on future data as reference, thereby serving as an indicator of its transferability across climates. A low value indicates that the model parameters are robust to climate shifts, whereas a high one suggests that a model trained on present-day data may not be directly applicable under future climate conditions.
However, the robustness gap defined in (12) is expressed in absolute terms, which makes its interpretation dependent on the scale of the loss function and of the variable under consideration. This lack of normalization complicates comparisons across different metrics, variables, or models. For instance, a robustness gap of 0.5 may indicate negligible degradation for one variable but severe deterioration for another, depending on the underlying scale. To address this limitation, we introduce normalized versions of the robustness gap that are more interpretable.
We consider two normalization strategies. Our first version of the normalized robustness gap is defined as
$\mathrm{NRG}_{\text{ratio}}=\frac{\left\langle\ell\left(y_{\text{F}},y_{\text{F}}^{\text{P}}\right)\right\rangle-\left\langle\ell\left(y_{\text{F}},y_{\text{F}}^{\text{F}}\right)\right\rangle}{\left\langle\ell\left(y_{\text{F}},y_{\text{F}}^{\text{F}}\right)\right\rangle}.$ (13)
Here, the denominator corresponds to the error of the model trained on future data, which serves as a theoretical benchmark. This formulation normalizes the average degradation by this benchmark error, providing an intuitive interpretation: a value of zero indicates no degradation, a value of one means the error doubles compared to the future-trained model, and values greater than one correspond to larger deterioration. Its main advantage lies in its simplicity and interpretability. However, this normalization can be sensitive to extreme pointwise benchmark errors because it relies on a single global mean, which is inherently non-robust to outliers.
Our second version of the normalized robustness gap is
$\mathrm{NRG}_{\text{pointwise}}=\left\langle\frac{\ell\left(y_{\text{F}},y_{\text{F}}^{\text{P}}\right)-\ell\left(y_{\text{F}},y_{\text{F}}^{\text{F}}\right)}{\ell\left(y_{\text{F}},y_{\text{F}}^{\text{F}}\right)}\right\rangle.$ (14)
In this case, the degradation relative to the future-trained model is computed for each grid point before averaging. This approach captures local variability and reduces the bias introduced by using a single global normalization factor, although it remains sensitive to extreme pointwise benchmark errors, which propagate into the expectation. Its interpretation is less straightforward than in the first case. In summary, the ratio of expectations in (13) emphasizes global interpretability while the expectation of pointwise ratios in (14) prioritizes statistical robustness.
All the metrics presented in this section (generalization gap, robustness gap, and normalized versions of the robustness gap) are computed for a specific dataset combining present-day and future conditions. They are inherently dataset-dependent and may differ across alternative datasets. To account for this variability, one could consider their expected values under the joint distribution of the variables appearing in the corresponding expressions (e.g., $y_{\text{F}},\,y_{\text{F}}^{\text{P}},\,y_{\text{F}}^{\text{F}}$ in the case of the robustness gap).
3.3.2 Special Case: Robustness Gap for Quantile Predictions
@@
Let $q_{\text{F},\alpha}^{\text{P}}$ denote the predicted quantile for the future climate obtained using a model trained on present-day data, and $q_{\text{F},\alpha}^{\text{F}}$ the predicted quantile obtained using a model trained on future data. The observed (true) quantiles in the present and future climate are denoted by $q_{\text{P},\alpha}$ and $q_{\text{F},\alpha}$, respectively. In this context, the robustness gap introduced above becomes
$\mathrm{RG}=\left\langle\ell(q_{\text{F},\alpha},q_{\text{F},\alpha}^{\text{P}})-\ell(q_{\text{F},\alpha},q_{\text{F},\alpha}^{\text{F}})\right\rangle.$ (15)
@@
We now introduce two interpretable quantities that will naturally appear in the decomposition of the pointwise robustness gap (see below). Let $\varepsilon_{\text{F},\alpha}=q_{\text{F},\alpha}^{\text{F}}-q_{\text{F},\alpha}$, which is the fit bias of the model trained on the future climate for the quantile at level $\alpha$, and $\Delta_{\alpha}=q_{\text{F},\alpha}^{\text{P}}-q_{\text{F},\alpha}^{\text{F}}$, as it quantifies the sensitivity of the predicted quantile to the choice of training data (whether from the present-day or future climate). We expect $\varepsilon_{F,\alpha}$ to be small if the model is adequate and has been properly trained. With these notations, it is straightforward to obtain that
$q_{\text{F},\alpha}^{\text{F}}=q_{\text{F},\alpha}+\varepsilon_{\text{F},\alpha},\quad q_{\text{F},\alpha}^{\text{P}}=q_{\text{F},\alpha}+\varepsilon_{\text{F},\alpha}+\Delta_{\alpha},$ (17)
and it can be shown that (see SM, Section A) that the pointwise robustness gap can be written
$\mathrm{PRG}$ $\overset{\mathrm{def}}{=}\ell_{\alpha}\left(q_{\text{F},\alpha},q^{\text{P}}_{\text{F},\alpha}\right)-\ell_{\alpha}\left(q_{\text{F},\alpha},q^{\text{F}}_{\text{F},\alpha}\right)$
@@
$\mathrm{PRG}=\begin{cases}(1-\alpha)\Delta_{\alpha},&\Delta_{\alpha}>0,\\ \alpha|\Delta_{\alpha}|,&\Delta_{\alpha}<0,\end{cases}$ (19)
@@
When incorporated into (3.3.2), this decomposition establishes a general framework for analyzing the contribution of individual parameter shifts to the pointwise robustness gap. It maintains the generality of the expression, ensuring applicability across a wide range of parametric distributions. Overall, our framework provides a comprehensive understanding of the robustness gap—clarifying not only when the super-resolution model generalizes effectively, when its performance begins to degrade, and when it ultimately fails, but also uncovering the underlying factors driving these behaviors.
@@ 4.1 Generalized Extreme-Value Distribution
To characterize extreme hourly precipitation, we adopt the Generalized Extreme-Value framework from Extreme-Value Theory (EVT), a branch of statistics focused on modeling the behavior of distribution tails. Unlike classical methods that describe central tendencies such as means and variances, EVT provides asymptotically justified models for block maxima or threshold exceedances, making it particularly well suited for assessing the risk of rare, high-impact events like extreme hourly rainfall (GEV).
@@ 4.2 Incorporating Features in the GEV Parameters
We now specify the form of the function $f$ introduced in (4), adapted to our specific setting in which the underlying parametric distribution is the GEV distribution. To flexibly capture how the features (parameters of the GEV distributions at coarse-resolution and auxiliary features) influence each parameter of the GEV distribution, we use the Vector Generalized Additive Model (VGAM) framework (Yee2015). The VGAM extends the familiar Generalized Additive Model (GAM) by allowing for multiple response variables. In our case the vector of response variables is composed of the GEV parameters.
@@
We typically model $\mu$ and $\log\sigma$ as smooth functions of the features but, for stability and identifiability, take $\xi$ to be constant (i.e., $f_{\mu,k}\equiv 0$ for all $k$). Indeed the estimation of $\xi$ is notoriously imprecise even with large samples of block maxima and allowing $\xi$ to vary with covariates often yields unstable fits (GEV; DavisonSmith1990). Moreover, extreme events are inherently rare, so there is typically insufficient data within each covariate “slice” to support a reliable smooth trend in $\xi$, and attempts to do so can lead to over‐parameterization and degraded predictive performance (EastoeTawn2009). Finally, empirical comparisons indicate that introducing time‐ or covariate‐dependence in the location (and sometimes scale) parameters captures the bulk of observed non‐stationarity, while varying $\xi$ provides only marginal improvements at the cost of substantially increased uncertainty (Katz2002; GEV).